OpenStudy (anonymous):

An alloy of gold and silver weighs 50 gms. It contains 80% gold. How much gold should be added to the alloy so that percentage of gold is increased to 90?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

can u plz provide the steps

OpenStudy (anonymous):

let x be the amount of gold that needs to be added: 40+x/50+x =0.9 45+0.9x=40+x 0.1x+5 x=50

OpenStudy (anonymous):

sorry 50 grams needs to be added

OpenStudy (king):

oh sry i misread the question!!!!!!!

OpenStudy (anonymous):

and obviously since the alloy is 40 gold 10 silver, adding 50 gold would make the total 90/100 gold which is 90%